# Test Time Learning for Time Series Forecasting

### v2

@@ 1 Introduction
Long Time Series Forecasting (LTSF) is a crucial task in various fields, including energy, industry, defense, and atmospheric sciences. LTSF uses a historical sequence of observations, known as the look-back window, to predict future values through a learned or mathematically induced model. However, the stochastic nature of real-world events makes LTSF challenging. Deep learning models, including time series forecasting, have been widely adopted in engineering and scientific fields. Early approaches employed Recurrent Neural Networks (RNNs) to capture long-range dependencies in sequential data like time series. However, recurrent architectures like RNNs have limited memory retention, are difficult to parallelize, and have constrained expressive capacity. Transformers [25], with ability to efficiently process sequential data in parallel and capture contextual information, have significantly improved performance on time series prediction task [26, 16, 20, 3]. Yet, due to the quadratic complexity of attention mechanisms with respect to the context window (or look-back window in LTSF), Transformers are limited in their ability to capture very long dependencies.
In recent years, State Space Models (SSMs) such as Mamba  [7], a gated linear RNN variant, have revitalized the use of RNNs for LTSF. These models efficiently capture much longer dependencies while reducing computational costs and enhancing expressive power and memory retention. A new class of Linear RNNs, known as Test Time Training (TTT) modules  [24], has emerged. These modules use expressive hidden states and provide theoretical guarantees for capturing long-range dependencies, positioning them as one of the most promising architectures for LTSF.
@@ 2 Related Work
Several Transformer-based models have advanced long-term time series forecasting (LTSF), with notable examples like iTransformer [18] and PatchTST
 [21]. iTransformer employs multimodal interactive attention to capture both temporal and inter-modal dependencies, suitable for multivariate time series, though it incurs high computational costs when multimodal data interactions are minimal. PatchTST, inspired by Vision Transformers [6], splits input sequences into patches to capture dependencies effectively, but its performance hinges on selecting the appropriate patch size and may reduce model interpretability. Other influential models include Informer [31], which uses sparse self-attention to reduce complexity but may overlook finer details in multivariate data; Autoformer [28], which excels in periodic data but struggles with non-periodic patterns; Pyraformer [16], which captures multi-scale dependencies through a hierarchical structure but at the cost of increased computational requirements; and Fedformer [32], which combines time- and frequency-domain representations for efficiency but may underperform on noisy time series. While each model advances LTSF in unique ways, they also introduce specific trade-offs and limitations.
@@ 3 Model Architecture
We incorporated TTT into the TimeMachine model, replacing the original Mamba block. We evaluated our approach with various setups, including different backbones and TTT layer configurations. Additionally, we introduced convolutional layers before the sequence modeling block and conducted experiments with different context lengths and prediction lengths.
@@ 3.1 General Architecture
Our model architecture builds upon the TimeMachine model [1], introducing key modifications, as shown in Figure 1, 1 and 1. Specifically, we replace the Mamba modules in TimeMachine with TTT (Test-Time Training) modules [24], which retain compatibility since both are linear RNNs [22]. However, TTT offers superior long-range dependency modeling due to its adaptive nature and theoretically infinite context window. A detailed visualization of the TTT block and the different proposed architectures can be found in Appendix B
Our model features a two-level hierarchical architecture that captures both high-resolution and low-resolution context, as illustrated in Figure 1. To adapts to the specific characteristics of the dataset, the architecture handles two scenarios—Channel Mixing and Channel Independence—illustrated in Figure 1 and 1 respectively. A more detailed and mathematical description of the normalization and prediction procedures can be found in Appendix B.
@@ 3.4 Final Prediction
After processing both high-resolution and low-resolution cues, the model concatenates the outputs from both levels. A final linear projection layer (P-2) is then applied to generate the output predictions. The output is subsequently passed through RevIN Denormalization, which reverses the initial normalization and maps the output back to its original scale for interpretability. For more detailed explanations and mathematical descriptions refer to Appendix B.
@@ 4.1 Original Experimental Setup
We evaluate our model on seven benchmark datasets that are commonly used for LTSF, namely: Traffic, Weather, Electricity, ETTh1, ETTh2, ETTm1, and ETTm2 from  [28] and  [31]. Among these, the Traffic and Electricity datasets are significantly larger, with 862 and 321 channels, respectively, and each containing tens of thousands of temporal points. Table 6 summarizes the dataset details in Appendix D.
For all experiments, we adopted the same setup as in [18], fixing the look-back window $L=96$ and testing four different prediction lengths $T={96,192,336,720}$. We compared our TimeMachine-TTT model against 12 state-of-the-art (SOTA) models, including TimeMachine  [1], iTransformer  [18], PatchTST  [21], DLinear  [29], RLinear  [14], Autoformer  [28], Crossformer  [30], TiDE  [4], Scinet  [15], TimesNet  [27], FEDformer  [32], and Stationary  [19]. All experiments were conducted with both MLP and Linear architectures using the original Mamba backbone, and we confirmed the results from the TimeMachine paper.
@@ 4.2 Quantitative Results
Across all seven benchmark datasets, our TimeMachine-TTT model consistently demonstrated superior performance compared to SOTA models. In the Weather dataset, TTT achieved leading performance at longer horizons (336 and 720), with MSEs of 0.248 and 0.341, respectively, outperforming TimeMachine, which recorded MSEs of 0.256 and 0.342. The Traffic dataset, with its high number of channels (862), also saw TTT outperform TimeMachine and iTransformer at both short (96-step MSE of 0.394 vs. 0.397) and long horizons (720-step MSE of 0.464 vs. 0.467), highlighting the model’s ability to handle multivariate time series data.
Table 1: Results in MSE and MAE (the lower the better) for the long-term forecasting task. We compare extensively with baselines under different prediction lengths, $T=\{96,192,336,720\}$ following the setting of iTransformer [17]. The length of the input sequence ($L$) is set to 96 for all baselines. TTT (ours) is our TTT block with the Conv Stack 5 architecture. The best results are in bold and the second best are underlined.
@@
Weather 96 0.163 0.211 0.164 0.208 0.174 0.214 0.192 0.232 0.177 0.218 0.158 0.230 0.202 0.261 0.172 0.220 0.196 0.255 0.221 0.306 0.217 0.296 0.173 0.223
192 0.220 0.258 0.211 0.250 0.221 0.254 0.240 0.271 0.225 0.259 0.206 0.277 0.242 0.298 0.219 0.261 0.237 0.296 0.261 0.340 0.276 0.336 0.245 0.285
336 0.248 0.279 0.256 0.290 0.278 0.296 0.292 0.307 0.278 0.297 0.272 0.335 0.287 0.335 0.280 0.306 0.283 0.335 0.309 0.378 0.339 0.380 0.321 0.338
720 0.341 0.342 0.342 0.343 0.358 0.349 0.364 0.353 0.354 0.348 0.398 0.418 0.351 0.386 0.365 0.359 0.345 0.381 0.377 0.427 0.403 0.428 0.414 0.410
Traffic 96 0.394 0.264 0.397 0.268 0.395 0.268 0.649 0.389 0.544 0.359 0.522 0.290 0.805 0.493 0.593 0.321 0.650 0.396 0.788 0.499 0.587 0.366 0.612 0.338
192 0.441 0.292 0.417 0.274 0.417 0.276 0.601 0.366 0.540 0.354 0.530 0.293 0.756 0.474 0.617 0.336 0.598 0.370 0.789 0.505 0.604 0.373 0.613 0.340
336 0.433 0.281 0.433 0.281 0.433 0.283 0.609 0.369 0.551 0.358 0.558 0.305 0.762 0.477 0.629 0.336 0.605 0.373 0.797 0.508 0.621 0.383 0.618 0.328
720 0.464 0.293 0.467 0.300 0.467 0.302 0.647 0.387 0.586 0.375 0.589 0.328 0.719 0.449 0.640 0.350 0.645 0.394 0.841 0.523 0.626 0.382 0.653 0.355
Electricity 96 0.140 0.233 0.142 0.236 0.148 0.240 0.201 0.281 0.195 0.285 0.219 0.314 0.237 0.329 0.168 0.272 0.197 0.282 0.247 0.345 0.193 0.308 0.169 0.273
192 0.158 0.256 0.158 0.250 0.162 0.253 0.201 0.283 0.199 0.289 0.231 0.322 0.236 0.330 0.184 0.289 0.196 0.285 0.257 0.355 0.201 0.315 0.182 0.286
336 0.169 0.256 0.172 0.268 0.178 0.269 0.215 0.298 0.215 0.305 0.246 0.337 0.249 0.344 0.198 0.300 0.209 0.301 0.269 0.369 0.214 0.329 0.200 0.304
720 0.201 0.290 0.207 0.298 0.225 0.317 0.257 0.331 0.256 0.337 0.280 0.363 0.284 0.373 0.220 0.320 0.245 0.333 0.299 0.390 0.246 0.355 0.222 0.321
ETTh1 96 0.358 0.379 0.364 0.387 0.386 0.405 0.386 0.395 0.414 0.419 0.423 0.448 0.479 0.464 0.384 0.402 0.386 0.400 0.654 0.599 0.376 0.419 0.513 0.491
192 0.409 0.415 0.415 0.416 0.441 0.436 0.437 0.424 0.460 0.445 0.471 0.474 0.525 0.492 0.436 0.429 0.437 0.432 0.719 0.631 0.420 0.448 0.534 0.504
@@
720 0.475 0.451 0.458 0.453 0.503 0.491 0.481 0.470 0.500 0.488 0.653 0.621 0.594 0.558 0.521 0.500 0.519 0.516 0.836 0.699 0.506 0.507 0.643 0.616
ETTh2 96 0.279 0.332 0.275 0.334 0.297 0.349 0.288 0.338 0.302 0.348 0.745 0.584 0.400 0.440 0.340 0.374 0.333 0.387 0.707 0.621 0.358 0.397 0.476 0.458
192 0.377 0.382 0.349 0.381 0.380 0.400 0.374 0.390 0.388 0.400 0.877 0.656 0.528 0.509 0.402 0.414 0.477 0.476 0.860 0.689 0.429 0.439 0.512 0.493
336 0.401 0.404 0.340 0.381 0.428 0.432 0.415 0.426 0.426 0.433 1.043 0.731 0.643 0.571 0.452 0.452 0.594 0.541 1.000 0.744 0.496 0.487 0.552 0.551
720 0.445 0.431 0.411 0.433 0.427 0.445 0.420 0.440 0.431 0.446 1.104 0.763 0.874 0.679 0.462 0.468 0.831 0.657 1.249 0.838 0.463 0.474 0.562 0.560
ETTm1 96 0.315 0.352 0.317 0.355 0.334 0.368 0.355 0.376 0.329 0.367 0.404 0.426 0.364 0.387 0.338 0.375 0.345 0.372 0.418 0.438 0.379 0.419 0.386 0.398
192 0.370 0.386 0.357 0.378 0.377 0.391 0.391 0.392 0.367 0.385 0.450 0.451 0.398 0.404 0.374 0.387 0.380 0.389 0.439 0.450 0.426 0.441 0.459 0.444
336 0.378 0.397 0.379 0.399 0.426 0.420 0.424 0.415 0.399 0.410 0.532 0.515 0.428 0.425 0.410 0.411 0.413 0.413 0.490 0.485 0.445 0.459 0.495 0.464
720 0.431 0.419 0.445 0.436 0.491 0.459 0.487 0.450 0.454 0.439 0.666 0.589 0.487 0.461 0.478 0.450 0.474 0.453 0.595 0.550 0.543 0.490 0.585 0.516
ETTm2 96 0.177 0.251 0.175 0.256 0.180 0.264 0.182 0.265 0.175 0.259 0.287 0.366 0.207 0.305 0.187 0.267 0.193 0.292 0.286 0.377 0.203 0.287 0.192 0.274
192 0.240 0.299 0.239 0.299 0.250 0.309 0.246 0.304 0.241 0.302 0.414 0.492 0.290 0.364 0.249 0.309 0.284 0.362 0.399 0.445 0.269 0.328 0.280 0.339
336 0.301 0.339 0.287 0.332 0.311 0.348 0.307 0.342 0.305 0.343 0.597 0.542 0.377 0.422 0.321 0.351 0.369 0.427 0.637 0.591 0.325 0.366 0.334 0.361
720 0.362 0.383 0.371 0.385 0.412 0.407 0.407 0.398 0.402 0.400 1.730 1.042 0.558 0.524 0.408 0.403 0.554 0.522 0.960 0.735 0.421 0.415 0.417 0.413
In the Electricity dataset, TTT showed dominant results across all horizons, achieving an MSE of 0.140 at horizon 96 and 0.201 at horizon 720, outperforming TimeMachine and PatchTST. For ETTh1, TTT was highly competitive, with strong short-term results (MSE of 0.358 at horizon 96) and continued dominance at medium-term horizons like 192, with an MSE of 0.409. While ETTh2 showed TimeMachine slightly ahead at short horizons, TTT closed the gap at longer horizons (MSE of 0.445 at horizon 720 compared to 0.411 for TimeMachine).
For the ETTm1 dataset, TTT outperformed TimeMachine at nearly every horizon, recording an MSE of 0.315 at horizon 96 and 0.431 at horizon 720, confirming its effectiveness for long-term industrial forecasting. Similarly, in ETTm2, TTT remained highly competitive at longer horizons, with a lead over TimeMachine at horizon 720 (MSE of 0.362 vs. 0.371). The radar plot in Figure 2 shows the comparison between TTT (ours) and TimeMachine for both MSE and MAE on all datasets.

### v3

@@ 1 Introduction
Long Time Series Forecasting (LTSF) is a crucial task in various fields, including energy [9], industry [10], defense [4], and atmospheric sciences [22]. LTSF uses a historical sequence of observations, known as the look-back window, to predict future values through a learned or mathematically induced model. However, the stochastic nature of real-world events makes LTSF challenging. Deep learning models, including time series forecasting, have been widely adopted in engineering and scientific fields. Early approaches employed Recurrent Neural Networks (RNNs) to capture long-range dependencies in sequential data like time series. However, recurrent architectures like RNNs have limited memory retention, are difficult to parallelize, and have constrained expressive capacity. Transformers [34], with ability to efficiently process sequential data in parallel and capture contextual information, have significantly improved performance on time series prediction task [36, 24, 28, 7]. Yet, due to the quadratic complexity of attention mechanisms with respect to the context window (or look-back window in LTSF), Transformers are limited in their ability to capture very long dependencies.
In recent years, State Space Models (SSMs) such as Mamba  [13], a gated linear RNN variant, have revitalized the use of RNNs for LTSF. These models efficiently capture much longer dependencies while reducing computational costs and enhancing expressive power and memory retention. A new class of Linear RNNs, known as Test Time Training (TTT) modules  [32], has emerged. These modules use expressive hidden states and provide theoretical guarantees for capturing long-range dependencies, positioning them as one of the most promising architectures for LTSF and due to their weight adaptation during test time are very effective on non-stationary data. We provide more motivation on TTT for non-stationary data in Appendix A.
@@ 2 Related Work
Several Transformer-based models have advanced long-term time series forecasting (LTSF), with notable examples like iTransformer [26] and PatchTST  [29]. iTransformer employs multimodal interactive attention to capture both temporal and inter-modal dependencies, suitable for multivariate time series, though it incurs high computational costs when multimodal data interactions are minimal. PatchTST, inspired by Vision Transformers [12], splits input sequences into patches to capture dependencies effectively, but its performance hinges on selecting the appropriate patch size and may reduce model interpretability. Other influential models include Informer [41], which uses sparse self-attention to reduce complexity but may overlook finer details in multivariate data; Autoformer [38], which excels in periodic data but struggles with non-periodic patterns; Pyraformer [24], which captures multi-scale dependencies through a hierarchical structure but at the cost of increased computational requirements; and Fedformer [42], which combines time- and frequency-domain representations for efficiency but may underperform on noisy time series. While each model advances LTSF in unique ways, they also introduce specific trade-offs and limitations.
@@ 3 Model Architecture
We incorporated TTT into the TimeMachine model, replacing the original Mamba block. We evaluated our approach with various setups, including different backbones and TTT layer configurations. Additionally, we introduced convolutional layers before the sequence modeling block and conducted experiments with different context lengths and prediction lengths. We provide mathematical foundations as to why TTT is able to perform test-time adaptation without catastrophic forgetting and how the module adapts to distribution shifts in Appendix A. In the same Appendix we quantify the computational overhead introduced by test-time updates and provide empirical validation, published by the authors who proposed TTT in [33], on how it performs on real corrupted data and provide some intuition on the parameter initialization in TTT as discussed in the same reference.
@@ 3.1 General Architecture
Our model architecture builds upon the TimeMachine model [1], introducing key modifications, as shown in Figure 1, 1 and 1. Specifically, we replace the Mamba modules in TimeMachine with TTT (Test-Time Training) modules [32], which retain compatibility since both are linear RNNs [30]. However, TTT offers superior long-range dependency modeling due to its adaptive nature and theoretically infinite context window. A detailed visualization of the TTT block and the different proposed architectures can be found in Appendix D
Our model features a two-level hierarchical architecture that captures both high-resolution and low-resolution context, as illustrated in Figure 1. To adapts to the specific characteristics of the dataset, the architecture handles two scenarios—Channel Mixing and Channel Independence—illustrated in Figure 1 and 1 respectively. A more detailed and mathematical description of the normalization and prediction procedures can be found in Appendix D. We provide a computational complexity analysis of the TTT, Transformer, Mamba and ModernTCN modules in Appendix F and we also provide a computational complexity analysis for our model, TimeMachine, iTransformer, PatchTST, TSMixer and ModernTCN in Appendix G. We also included a mathematical comparison between the Mamba and TTT modules in Appendix B as well as theoretical comparison between the TTT module and models handling noise or temporal regularization in Appendix C.
@@ 3.2 Hierarchical Embedding
We provide more intuition on the effectivenes of hierarchical modeling in Appendix E.
@@ 3.4 Final Prediction
After processing both high-resolution and low-resolution cues, the model concatenates the outputs from both levels. A final linear projection layer (P-2) is then applied to generate the output predictions. The output is subsequently passed through RevIN Denormalization, which reverses the initial normalization and maps the output back to its original scale for interpretability. For more detailed explanations and mathematical descriptions refer to Appendix D.
@@ 4.1 Original Experimental Setup
We evaluate our model on seven benchmark datasets that are commonly used for LTSF, namely: Traffic, Weather, Electricity, ETTh1, ETTh2, ETTm1, and ETTm2 from  [38] and  [41]. Among these, the Traffic and Electricity datasets are significantly larger, with 862 and 321 channels, respectively, and each containing tens of thousands of temporal points. Table 6 summarizes the dataset details in Appendix I.
For all experiments, we adopted the same setup as in [26], fixing the look-back window $L=96$ and testing four different prediction lengths $T={96,192,336,720}$. We compared our TimeMachine-TTT model against 12 state-of-the-art (SOTA) models, including TimeMachine  [1], iTransformer  [26], PatchTST  [29], DLinear  [39], RLinear  [21], Autoformer  [38], Crossformer  [40], TiDE  [8], Scinet  [23], TimesNet  [37], FEDformer  [42], and Stationary  [27]. All experiments were conducted with both MLP and Linear architectures using the original Mamba backbone, and we confirmed the results from the TimeMachine paper. We include calculations on the resource utilization of the model in Appendix G and quantify the impact of test-time updates on memory and latency in Appendix A.
@@ 4.2 Quantitative Results
Across all seven benchmark datasets, our TimeMachine-TTT model consistently demonstrated superior performance compared to SOTA models. In the Weather dataset, TTT achieved leading performance at longer horizons (336 and 720), with MSEs of 0.246 and 0.339, respectively, outperforming TimeMachine, which recorded MSEs of 0.256 and 0.342. The Traffic dataset, with its high number of channels (862), also saw TTT outperform TimeMachine and iTransformer at both medium (336-step MSE of 0.430 vs. 0.433) and long horizons (720-step MSE of 0.464 vs. 0.467), highlighting the model’s ability to handle multivariate time series data.
Table 1: Results in MSE and MAE (the lower the better) for the long-term forecasting task (averaged over 5 runs). We compare extensively with baselines under different prediction lengths, $T=\{96,192,336,720\}$ following the setting of iTransformer [25]. The length of the input sequence ($L$) is set to 96 for all baselines. TTT (ours) is our TTT block with the Conv Stack 5 architecture. The best results are in bold and the second best are underlined.
@@
Weather 96 0.165 0.214 0.164 0.208 0.174 0.214 0.192 0.232 0.177 0.218 0.158 0.230 0.202 0.261 0.172 0.220 0.196 0.255 0.221 0.306 0.217 0.296 0.173 0.223
192 0.225 0.263 0.211 0.250 0.221 0.254 0.240 0.271 0.225 0.259 0.206 0.277 0.242 0.298 0.219 0.261 0.237 0.296 0.261 0.340 0.276 0.336 0.245 0.285
336 0.246 0.275 0.256 0.290 0.278 0.296 0.292 0.307 0.278 0.297 0.272 0.335 0.287 0.335 0.280 0.306 0.283 0.335 0.309 0.378 0.339 0.380 0.321 0.338
720 0.339 0.343 0.342 0.343 0.358 0.349 0.364 0.353 0.354 0.348 0.398 0.418 0.351 0.386 0.365 0.359 0.345 0.381 0.377 0.427 0.403 0.428 0.414 0.410
Traffic 96 0.397 0.268 0.397 0.268 0.395 0.268 0.649 0.389 0.544 0.359 0.522 0.290 0.805 0.493 0.593 0.321 0.650 0.396 0.788 0.499 0.587 0.366 0.612 0.338
192 0.434 0.287 0.417 0.274 0.417 0.276 0.601 0.366 0.540 0.354 0.530 0.293 0.756 0.474 0.617 0.336 0.598 0.370 0.789 0.505 0.604 0.373 0.613 0.340
336 0.430 0.283 0.433 0.281 0.433 0.283 0.609 0.369 0.551 0.358 0.558 0.305 0.762 0.477 0.629 0.336 0.605 0.373 0.797 0.508 0.621 0.383 0.618 0.328
720 0.456 0.286 0.467 0.300 0.467 0.302 0.647 0.387 0.586 0.375 0.589 0.328 0.719 0.449 0.640 0.350 0.645 0.394 0.841 0.523 0.626 0.382 0.653 0.355
Electricity 96 0.135 0.230 0.142 0.236 0.148 0.240 0.201 0.281 0.195 0.285 0.219 0.314 0.237 0.329 0.168 0.272 0.197 0.282 0.247 0.345 0.193 0.308 0.169 0.273
192 0.153 0.254 0.158 0.250 0.162 0.253 0.201 0.283 0.199 0.289 0.231 0.322 0.236 0.330 0.184 0.289 0.196 0.285 0.257 0.355 0.201 0.315 0.182 0.286
336 0.166 0.255 0.172 0.268 0.178 0.269 0.215 0.298 0.215 0.305 0.246 0.337 0.249 0.344 0.198 0.300 0.209 0.301 0.269 0.369 0.214 0.329 0.200 0.304
720 0.199 0.285 0.207 0.298 0.225 0.317 0.257 0.331 0.256 0.337 0.280 0.363 0.284 0.373 0.220 0.320 0.245 0.333 0.299 0.390 0.246 0.355 0.222 0.321
ETTh1 96 0.352 0.375 0.364 0.387 0.386 0.405 0.386 0.395 0.414 0.419 0.423 0.448 0.479 0.464 0.384 0.402 0.386 0.400 0.654 0.599 0.376 0.419 0.513 0.491
192 0.412 0.418 0.415 0.416 0.441 0.436 0.437 0.424 0.460 0.445 0.471 0.474 0.525 0.492 0.436 0.429 0.437 0.432 0.719 0.631 0.420 0.448 0.534 0.504
@@
720 0.478 0.454 0.458 0.453 0.503 0.491 0.481 0.470 0.500 0.488 0.653 0.621 0.594 0.558 0.521 0.500 0.519 0.516 0.836 0.699 0.506 0.507 0.643 0.616
ETTh2 96 0.274 0.328 0.275 0.334 0.297 0.349 0.288 0.338 0.302 0.348 0.745 0.584 0.400 0.440 0.340 0.374 0.333 0.387 0.707 0.621 0.358 0.397 0.476 0.458
192 0.373 0.379 0.349 0.381 0.380 0.400 0.374 0.390 0.388 0.400 0.877 0.656 0.528 0.509 0.402 0.414 0.477 0.476 0.860 0.689 0.429 0.439 0.512 0.493
336 0.403 0.408 0.340 0.381 0.428 0.432 0.415 0.426 0.426 0.433 1.043 0.731 0.643 0.571 0.452 0.452 0.594 0.541 1.000 0.744 0.496 0.487 0.552 0.551
720 0.448 0.434 0.411 0.433 0.427 0.445 0.420 0.440 0.431 0.446 1.104 0.763 0.874 0.679 0.462 0.468 0.831 0.657 1.249 0.838 0.463 0.474 0.562 0.560
ETTm1 96 0.309 0.348 0.317 0.355 0.334 0.368 0.355 0.376 0.329 0.367 0.404 0.426 0.364 0.387 0.338 0.375 0.345 0.372 0.418 0.438 0.379 0.419 0.386 0.398
192 0.371 0.389 0.357 0.378 0.377 0.391 0.391 0.392 0.367 0.385 0.450 0.451 0.398 0.404 0.374 0.387 0.380 0.389 0.439 0.450 0.426 0.441 0.459 0.444
336 0.381 0.401 0.379 0.399 0.426 0.420 0.424 0.415 0.399 0.410 0.532 0.515 0.428 0.425 0.410 0.411 0.413 0.413 0.490 0.485 0.445 0.459 0.495 0.464
720 0.433 0.423 0.445 0.436 0.491 0.459 0.487 0.450 0.454 0.439 0.666 0.589 0.487 0.461 0.478 0.450 0.474 0.453 0.595 0.550 0.543 0.490 0.585 0.516
ETTm2 96 0.180 0.253 0.175 0.256 0.180 0.264 0.182 0.265 0.175 0.259 0.287 0.366 0.207 0.305 0.187 0.267 0.193 0.292 0.286 0.377 0.203 0.287 0.192 0.274
192 0.242 0.301 0.239 0.299 0.250 0.309 0.246 0.304 0.241 0.302 0.414 0.492 0.290 0.364 0.249 0.309 0.284 0.362 0.399 0.445 0.269 0.328 0.280 0.339
336 0.302 0.341 0.287 0.332 0.311 0.348 0.307 0.342 0.305 0.343 0.597 0.542 0.377 0.422 0.321 0.351 0.369 0.427 0.637 0.591 0.325 0.366 0.334 0.361
720 0.364 0.384 0.371 0.385 0.412 0.407 0.407 0.398 0.402 0.400 1.730 1.042 0.558 0.524 0.408 0.403 0.554 0.522 0.960 0.735 0.421 0.415 0.417 0.413
In the Electricity dataset, TTT showed dominant results across all horizons, achieving an MSE of 0.135, 0.153, 0.166 and 0.199 at horizons 96, 192, 336, and 720 respectively, outperforming TimeMachine and PatchTST. For ETTh1, TTT was highly competitive, with strong short-term results (MSE of 0.352 at horizon 96) and continued dominance at medium-term horizons like 336, with an MSE of 0.412. For ETTh2, TTT beat TimeMachine on horizon 96 (MSE of 0.274), TTT also closed the gap at longer horizons (MSE of 0.448 at horizon 720 compared to 0.411 for TimeMachine).
For the ETTm1 dataset, TTT outperformed TimeMachine at nearly every horizon, recording an MSE of 0.309, 0.381 and 0.431 at horizon 96, 336 and 720 respectively, confirming its effectiveness for long-term forecasting. Similarly, in ETTm2, TTT remained highly competitive at longer horizons, with a lead over TimeMachine at horizon 720 (MSE of 0.362 vs. 0.371). The radar plot in Figure 2 shows the comparison between TTT (ours) and TimeMachine for both MSE and MAE on all datasets.